OpenStudy (anonymous):

simplify 3m-9m^2/-3m

OpenStudy (anonymous):

3m-9m^2 /-3m \[\left( 3m-9m ^{2} \right)/-3m\] \[\frac{ 3m \left( 1-3m \right) }{ -3m }\] \[=-\left( 1-3m \right)=-1+3m\]